# RoboRefer: Towards Spatial Referring with Reasoning in Vision-Language Models for Robotics

### v3

@@ 3.3.2 Data Recipe
In Fig. 3, we present the dataset recipe that progressively integrates 2D, 3D, and simulated data to enable general VLMs to adapt to spatial referring tasks, thereby enhancing bottom-up spatial understanding and reasoning. (1) 2D Web Images aim to endow the model with core spatial concepts and comprehensive depth perception across both indoor and outdoor scenes. To mitigate depth scale and category discrepancies between indoor and outdoor scenes, we leverage the large-scale, diverse 2D web image dataset, OpenImage [132]. However, directly extracting 3D-aware spatial information is challenging. Inspired by prior work [1, 35], we transform 2D images into pseudo-3D scene graphs. In detail, after high-quality filtering (from 1.7M to 466K images), we further enhance the data using Qwen2.5-VL [11] and a heuristic for generating hierarchical region captions, capturing both coarse labels and fine-grained spatial references, differentiating our approach from previous methods. We then construct scene graphs via object detection/segmentation, depth estimation, and camera intrinsic estimation, using object captions as nodes and spatial relations as edges. Finally, we generate QA pairs via template-based or LLM-based approaches, augmented by object-location QA derived from the annotated captions. (2) 3D Embodied Videos want to provide the model with a focused spatial understanding of indoor scenes, with a finer-grained perception of spatial relations and concepts. We therefore leverage the richly annotated CA-1M [133]. After fine-grained filtering (from 2M to 100K frames), we construct 3D scene graphs with more diverse spatial relations, enabled by precise 3D bounding boxes compared to 2D approaches. Moreover, we generate top-down occupancy maps that encode the object positions, orientations, and metric distances (e.g., “10cm right of the chair”), enabling accurate spatial referring for placement. (3) Simulation Data arms the model with multi-step referring capabilities with spatial reasoning. While 2D and 3D data enable single-step spatial understanding, they are less scalable for multi-step spatial referring with reasoning. Therefore, we leverage procedurally generate scene layouts [134], using manually verified assets [135] (from 46k to 3k) with semantic orientation annotations [7]. Tasks are purposefully designed to foster multi-step spatial referring and generate corresponding data. We assume that the generated code reflects optimal reasoning, with each line translated into textual form and intermediate results filled into structured formats (e.g., coordinates, distances), as shown in Fig. 2, Fig. 3, and Appx. D.4.2, yielding QA pairs with reasoning annotations. For more demonstrations about RefSpatial, please refer to Appx. F.
@@ 3.4 Training Details
We adopt NVILA [37] (2B/8B) as the base model and apply SFT to obtain RoboRefer-SFT. Due to computational limits, RFT is applied only to the 2B model, yielding RoboRefer-RFT. SFT has two steps: the first uses only the RefSpatial; the second trains on a mixture of RefSpatial, instruction tuning (1/20 the size of RefSpatial QA)[130, 131], and referring datasets[86]. Notably, RefSpatial is reused with both RGB and RGB-D inputs in the second step to enforce the image encoder to learn spatial understanding beyond depth cues. Thus, the model supports both RGB-only and RGB-D inference, with depth optionally inferred via a relative depth estimation model [136]. Finally, RFT stage uses the multi-step reasoning data from RefSpatial to train. See Appx. D for details.
@@ 4 Experiments
Method Input CV-Bench [15] $\text{BLINK}_{val}$ [16] RoboSpatial [2] SAT [4] EmbSpatial [22]
@@
GPT-4o [8] RGB 84.62 86.50 83.33 82.52 78.23 77.20 68.67 63.38
Gemini-2.5-Pro [9] RGB 93.54 91.00 90.67 91.61 87.90 77.24 70.59 76.67
Claude-3.7-Sonnet [36] RGB 74.15 85.83 84.17 74.83 67.74 60.73 40.67 33.33
@@
NVILA-2B [37] RGB 70.15 79.67 60.00 67.83 62.10 51.79 31.33 47.34
NVILA-8B [37] RGB 91.54 91.83 90.67 76.92 76.61 59.35 63.33 67.72
Qwen-2.5-VL-7B [11] RGB 82.15 60.17 69.00 64.34 60.98 49.59 30.00 40.20
Qwen-2.5-VL-72B [11] RGB 84.15 86.17 84.15 78.32 73.55 70.73 65.33 57.69
@@
SpatialBot-3B [14] RGB-D 69.38 77.33 60.83 67.83 67.74 72.36 63.33 50.66
SpatialRGPT-8B [1] RGB-D 91.00 89.8 88.50 81.12 89.51 66.67 64.00 59.62
SpaceLLaVA-13B [6] RGB 63.69 66.83 70.17 72.73 62.90 61.00 62.67 49.40
RoboPoint-13B [5] RGB 75.85 77.83 44.50 60.84 61.29 69.90 46.60 49.31
@@
Gemini-2.5-Pro [9] SpaceLLaVA [6] RoboPoint [5] Molmo-7B [15] Molmo-72B [15] 2B-SFT 8B-SFT 2B-RFT
@@ 4.1 Single-step Spatial Understanding
We evaluate on public single-step spatial understanding benchmarks, including CV-Bench [15], the BLINK [16] validation split, RoboSpatial [2] configuration part, SAT [4], and EmbSpatial [22]. Check Appx. E.2 for more evaluation details. The following parts present our analyses.
@@ 4.2 Multi-step Spatial Referring
We first evaluate current robotic referring benchmarks, namely RoboRefIt [137] (location) and Where2Place [5]/RoboSpatial [2] (placement), all limited to $2$ reasoning steps. To evaluate more complex multi-step spatial referring, we propose RefSpatial-Bench, a challenging benchmark based on real-world cluttered scenes. It contains two subsets, Location and Placement, each with $100$ images. Notably, $77$ images involve spatial relation combinations unseen in RefSpatial. Over $70\%$ requires multi-step reasoning (up to $5$ steps), including precise ground-truth masks. More details about RefSpatial can be found in Appx. C. For metrics, we report the average success rate of predicted points within the mask. We evaluate RoboRefer using RGB-D inputs by default, with depth maps generated from RGB images via DepthAnything V2 [136]. See Appx. E.3 for more details.
@@
Benchmark NVILA-2B [37] RoboRefer-2B-SFT
@@
Figure 4: RefSpatial-Bench results. G.P., $M.$, and R.P. donate Gemini-2.5-Pro [9], Molmo-72B [15], and RoboPoint [5]. RoboRefer-RFT excels in unseen and multi-step cases.
RFT stage fosters better reasoning ability. As shown in Tab. 2, the 2B-RFT variant outperforms all baselines, exceeding the prior SOTA (Gemini-2.5-Pro [9]) by 17.4% (absolute) on RefSpatial-Bench. We find that although Gemini-2.5-Pro excels in 2D referring (e.g., color, image-space localization), it struggles with 3D spatial relations involving distance (e.g., identifying the second-farthest object), reducing overall performance when multiple spatial constraints are combined. Fig. 4 shows complex multi-step spatial reasoning cases from RefSpatial-Bench with model comparisons.
RFT stage provides powerful generalization ability. In the RefSpatial-Bench-Unseen row of Tab. 2, we evaluate novel spatial relation combinations omitted during our SFT/RFT training. The 2B-RFT model exceeds 2B-SFT by 9.1% (absolute), indicating that SFT overfits the training distribution, whereas the RFT model better generalizes by leveraging learned spatial knowledge, consistent with prior findings [138]. Fig. 4 shows results on these unseen combinations across different models.
@@ 4.3 Public vision-language benchmarks
RefSpatial enhance 2D general referring ability. We also evaluate 2D referring capability on the ReCOCO/+/g [86]. Since our model predicts a single point, we deem a prediction correct if the point lies within the ground-truth bounding box. As this evaluation differs from standard visual grounding protocols, we additionally assess Qwen-2.5VL-72B [11] by using either its predicted point or the center of its predicted box as baselines. In Tab. 4, our method surpasses baselines, indicating that our dataset not only supports 3D spatial referring but also enhances 2D referring performance.
Joint RGB and RGB-D training preserves commonsense knowledge. In Tab. 4, we assess how spatial and depth information influences overall VQA performance by comparing RoboRefer-2B-SFT with the baseline NVILA-2B [37], trained on standard VQA datasets. Our model achieves comparable or slightly superior results, corroborating insights from SpatialVLM [6] and SpatialRGPT [1]. These findings indicate that although VLMs often struggle with spatial reasoning, targeted spatial VQA training, especially with combined RGB and RGB-D data enriched by general visual instruction datasets, can enhance spatial understanding without compromising overall VQA performance.
@@ 4.4 Simulator and Real-world Evaluation for Robotics
RoboRefer can be integrated into the system as a useful tool. We evaluate our model on the Open6DOR [58] V2 position track, comparing against VLA-based baselines (pretrained Octo [139], LIBERO-finetuned OpenVLA [140]) and SoFar [7], which integrates Florence-2 [92], SAM [77], GPT-4o, and GSNet. RoboRefer serves as a lightweight alternative to Florence-2 and GPT-4o for object localization and placement. By using a single target point predicted by RoboRefer, the system can generate more accurate masks and corresponding grasp poses than those from 2D boxes under occlusion in cluttered scenes, yielding a 6.8% absolute improvement in success rate (Tab. 6). Its compact size also reduces execution time by 27.5% relative to GPT-4o. See Appx. E.4 for details.
@@ 4.5 Ablation Study
Data recipe is critical for SFT training. Ablation results in Tab. 7 reveal that combining 2D, 3D, and simulated data yields optimal performance. As noted in Sec. 3.3, 2D data spans indoor/outdoor scenes, enabling depth learning across scales; its removal severely degrades performance on outdoor-centric BLINK [16]. Meanwhile, 3D data captures embodied indoor environments and mitigates the Sim2Real gap, benefiting indoor-focused CV-Bench [15]. Finally, simulated data broadens spatial diversity. This tripartite data composition is thus key to effective SFT training.
Dedicated depth encoder preserves image understanding. We compare dedicated and shared image-depth encoders during SFT. In Tab. 4, the dedicated encoder better maintains image understanding under limited RGB-only data (1/20 RefSpatial QA), while the shared encoder harms general performance. Though prior work[1] adopts a shared encoder, it (1) requires over twice as much RGB-only data compared to spatial-related data for co-training; (2) targets region-level depths, differing from our full-image approach.
Depth encoder improves both spatial understanding and reasoning. Recent VLMs [106, 3] show that large-scale spatial training enables implicit 3D understanding (e.g., depth, distance, 3D boxes) from images alone. To assess this, we fine-tune NVILA-2B [37] on RefSpatial without the depth encoder, followed by continued RFT. Results indicate that depth improves single-step spatial understanding, consistent with MM-Spatial [3], and yields greater gains in multi-step spatial referring. We attribute this to: (1) the need for precise coordinate prediction in spatial referring, unlike VQA’s multiple-choice; (2) cumulative reasoning across steps, amplifying the utility of depth cues.
@@ Appendix A Discussion
Depth-to-3D mapping assumption. The depth-to-3D mapping assumption is essential in our real-world evaluation, as our model predicts only 2D image-plane points, while real-world tasks typically require 3D coordinates for grasping, placement, or navigation. While depth noise and partial observations are important real-world challenges, our setting follows prior work [44, 5], which assumes that accurate depth-to-3D mapping is feasible given known camera intrinsics and extrinsics—sufficient for common manipulation and navigation tasks. Moreover, these challenges can be effectively mitigated via the following strategies: (1) Depth noise can be mitigated by recent advances in monocular depth estimation [67], monocular geometry prediction [141], and stereo methods [142]. In cases of severe noise, we employ FoundationStereo [142] in real-world settings to mitigate this issue. (2) Partial views are mitigated in our method by leveraging pixel-level target points. Further improvement is possible by incorporating RoboRefer as a spatially-aware planner for active perception.
@@
Method CV-Bench [15] $\text{BLINK}_{val}$ [16] RoboSpatial [2] SAT [4] EmbSpatial [22]
@@ Appendix B Implementation Details and Samples of RefSpatial Dataset
3D Embodied Video (Appx. B.2): This section outlines the 3D data selection process from CA-1M [133], discusses its limitations and mitigation strategies, and presents methods for enriched scene graph construction compared to the 2D data source. We further describe a QA generation framework that leverages detailed 3D annotations (e.g., depth maps, oriented 3D bounding boxes) to capture richer spatial relations. Finally, we detail how to generate QA pairs for the problem of “feasibility assessment for object placement in free space”.

### v4

@@ 3.3.2 Data Recipe
In Fig. 3, we present the dataset recipe that progressively integrates 2D, 3D, and simulated data to enable general VLMs to adapt to spatial referring tasks, thereby enhancing bottom-up spatial understanding and reasoning. (1) 2D Web Images aim to endow the model with core spatial concepts and comprehensive depth perception across both indoor and outdoor scenes. To mitigate depth scale and category discrepancies between indoor and outdoor scenes, we leverage the large-scale, diverse 2D web image dataset, OpenImage [kuznetsova2020open]. However, directly extracting 3D-aware spatial information is challenging. Inspired by prior work [cheng2024spatialrgpt, ma2025spatialreasoner], we transform 2D images into pseudo-3D scene graphs. In detail, after high-quality filtering (from 1.7M to 466K images), we further enhance the data using Qwen2.5-VL [bai2025qwen2] and a heuristic for generating hierarchical region captions, capturing both coarse labels and fine-grained spatial references, differentiating our approach from previous methods. We then construct scene graphs via object detection/segmentation, depth estimation, and camera intrinsic estimation, using object captions as nodes and spatial relations as edges. Finally, we generate QA pairs via template-based or LLM-based approaches, augmented by object-location QA derived from the annotated captions. (2) 3D Embodied Videos want to provide the model with a focused spatial understanding of indoor scenes, with a finer-grained perception of spatial relations and concepts. We therefore leverage the richly annotated CA-1M [lazarow2024cubify]. After fine-grained filtering (from 2M to 100K frames), we construct 3D scene graphs with more diverse spatial relations, enabled by precise 3D bounding boxes compared to 2D approaches. Moreover, we generate top-down occupancy maps that encode the object positions, orientations, and metric distances (e.g., “10cm right of the chair”), enabling accurate spatial referring for placement. (3) Simulation Data arms the model with multi-step referring capabilities with spatial reasoning. While 2D and 3D data enable single-step spatial understanding, they are less scalable for multi-step spatial referring with reasoning. Therefore, we leverage procedurally generate scene layouts [raistrick2024infinigen], using manually verified assets [deitke2023objaverse] (from 46k to 3k) with semantic orientation annotations [qi2025sofar]. Tasks are purposefully designed to foster multi-step spatial referring and generate corresponding data. We assume that the generated code reflects optimal reasoning, with each line translated into textual form and intermediate results filled into structured formats (e.g., coordinates, distances), as shown in Fig. 2, Fig. 3, and Appx. D.4.2, yielding QA pairs with reasoning annotations. For more demonstrations about RefSpatial, please refer to Appx. F.
@@ 3.4 Training Details
We adopt NVILA [liu2024nvila] (2B/8B) as the base model and apply SFT to obtain RoboRefer-SFT. Due to computational limits, RFT is applied only to the 2B model, yielding RoboRefer-RFT. SFT has two steps: the first uses only the RefSpatial; the second trains on a mixture of RefSpatial, instruction tuning (1/20 the size of RefSpatial QA)[liu2024mitigating, liu2024improved], and referring datasets[yu2016modeling]. Notably, RefSpatial is reused with both RGB and RGB-D inputs in the second step to enforce the image encoder to learn spatial understanding beyond depth cues. Thus, the model supports both RGB-only and RGB-D inference, with depth optionally inferred via a relative depth estimation model [yang2024depth]. Finally, RFT stage uses the multi-step reasoning data from RefSpatial to train. See Appx. D for details.
@@ 4 Experiments
Method Input CV-Bench [tong2024cambrian] $\text{BLINK}_{val}$ [fu2024blink] RoboSpatial [song2024robospatial] SAT [ray2024sat] EmbSpatial [du2024embspatial]
@@
GPT-4o [achiam2023gpt] RGB 84.62 86.50 83.33 82.52 78.23 77.20 68.67 63.38
Gemini-2.5-Pro [team2023gemini] RGB 93.54 91.00 90.67 91.61 87.90 77.24 70.59 76.67
Claude-3.7-Sonnet [anthropic2024claude] RGB 74.15 85.83 84.17 74.83 67.74 60.73 40.67 33.33
@@
NVILA-2B [liu2024nvila] RGB 70.15 79.67 60.00 67.83 62.10 51.79 31.33 47.34
NVILA-8B [liu2024nvila] RGB 91.54 91.83 90.67 76.92 76.61 59.35 63.33 67.72
Qwen-2.5-VL-7B [bai2025qwen2] RGB 82.15 60.17 69.00 64.34 60.98 49.59 30.00 40.20
Qwen-2.5-VL-72B [bai2025qwen2] RGB 84.15 86.17 84.15 78.32 73.55 70.73 65.33 57.69
@@
SpatialBot-3B [cai2024spatialbot] RGB-D 69.38 77.33 60.83 67.83 67.74 72.36 63.33 50.66
SpatialRGPT-8B [cheng2024spatialrgpt] RGB-D 91.00 89.8 88.50 81.12 89.51 66.67 64.00 59.62
SpaceLLaVA-13B [chen2024spatialvlm] RGB 63.69 66.83 70.17 72.73 62.90 61.00 62.67 49.40
RoboPoint-13B [yuan2024robopoint] RGB 75.85 77.83 44.50 60.84 61.29 69.90 46.60 49.31
@@
Gemini-2.5-Pro [team2023gemini] SpaceLLaVA [chen2024spatialvlm] RoboPoint [yuan2024robopoint] Molmo-7B [tong2024cambrian] Molmo-72B [tong2024cambrian] 2B-SFT 8B-SFT 2B-RFT
@@ 4.1 Single-step Spatial Understanding
We evaluate on public single-step spatial understanding benchmarks, including CV-Bench [tong2024cambrian], the BLINK [fu2024blink] validation split, RoboSpatial [song2024robospatial] configuration part, SAT [ray2024sat], and EmbSpatial [du2024embspatial]. Check Appx. E.2 for more evaluation details. The following parts present our analyses.
@@ 4.2 Multi-step Spatial Referring
We first evaluate current robotic referring benchmarks, namely RoboRefIt [lu2023vl] (location) and Where2Place [yuan2024robopoint]/RoboSpatial [song2024robospatial] (placement), all limited to $2$ reasoning steps. To evaluate more complex multi-step spatial referring, we propose RefSpatial-Bench, a challenging benchmark based on real-world cluttered scenes. It contains two subsets, Location and Placement, each with $100$ images. Notably, $77$ images involve spatial relation combinations unseen in RefSpatial. Over $70\%$ requires multi-step reasoning (up to $5$ steps), including precise ground-truth masks. More details about RefSpatial can be found in Appx. C. For metrics, we report the average success rate of predicted points within the mask. We evaluate RoboRefer using RGB-D inputs by default, with depth maps generated from RGB images via DepthAnything V2 [yang2024depth]. See Appx. E.3 for more details.
@@
Benchmark NVILA-2B [liu2024nvila] RoboRefer-2B-SFT
@@
Figure 4: RefSpatial-Bench results. G.P., $M.$, and R.P. donate Gemini-2.5-Pro [team2023gemini], Molmo-72B [tong2024cambrian], and RoboPoint [yuan2024robopoint]. RoboRefer-RFT excels in unseen and multi-step cases.
RFT stage fosters better reasoning ability. As shown in Tab. 2, the 2B-RFT variant outperforms all baselines, exceeding the prior SOTA (Gemini-2.5-Pro [team2023gemini]) by 17.4% (absolute) on RefSpatial-Bench. We find that although Gemini-2.5-Pro excels in 2D referring (e.g., color, image-space localization), it struggles with 3D spatial relations involving distance (e.g., identifying the second-farthest object), reducing overall performance when multiple spatial constraints are combined. Fig. 4 shows complex multi-step spatial reasoning cases from RefSpatial-Bench with model comparisons.
RFT stage provides powerful generalization ability. In the RefSpatial-Bench-Unseen row of Tab. 2, we evaluate novel spatial relation combinations omitted during our SFT/RFT training. The 2B-RFT model exceeds 2B-SFT by 9.1% (absolute), indicating that SFT overfits the training distribution, whereas the RFT model better generalizes by leveraging learned spatial knowledge, consistent with prior findings [chu2025sft]. Fig. 4 shows results on these unseen combinations across different models.
@@ 4.3 Public vision-language benchmarks
RefSpatial enhance 2D general referring ability. We also evaluate 2D referring capability on the ReCOCO/+/g [yu2016modeling]. Since our model predicts a single point, we deem a prediction correct if the point lies within the ground-truth bounding box. As this evaluation differs from standard visual grounding protocols, we additionally assess Qwen-2.5VL-72B [bai2025qwen2] by using either its predicted point or the center of its predicted box as baselines. In Tab. 4, our method surpasses baselines, indicating that our dataset not only supports 3D spatial referring but also enhances 2D referring performance.
Joint RGB and RGB-D training preserves commonsense knowledge. In Tab. 4, we assess how spatial and depth information influences overall VQA performance by comparing RoboRefer-2B-SFT with the baseline NVILA-2B [liu2024nvila], trained on standard VQA datasets. Our model achieves comparable or slightly superior results, corroborating insights from SpatialVLM [chen2024spatialvlm] and SpatialRGPT [cheng2024spatialrgpt]. These findings indicate that although VLMs often struggle with spatial reasoning, targeted spatial VQA training, especially with combined RGB and RGB-D data enriched by general visual instruction datasets, can enhance spatial understanding without compromising overall VQA performance.
@@ 4.4 Simulator and Real-world Evaluation for Robotics
RoboRefer can be integrated into the system as a useful tool. We evaluate our model on the Open6DOR [ding2024open6dor] V2 position track, comparing against VLA-based baselines (pretrained Octo [team2024octo], LIBERO-finetuned OpenVLA [kimopenvla]) and SoFar [qi2025sofar], which integrates Florence-2 [xiao2024florence], SAM [ravi2024sam], GPT-4o, and GSNet. RoboRefer serves as a lightweight alternative to Florence-2 and GPT-4o for object localization and placement. By using a single target point predicted by RoboRefer, the system can generate more accurate masks and corresponding grasp poses than those from 2D boxes under occlusion in cluttered scenes, yielding a 6.8% absolute improvement in success rate (Tab. 6). Its compact size also reduces execution time by 27.5% relative to GPT-4o. See Appx. E.4 for details.
@@ 4.5 Ablation Study
Data recipe is critical for SFT training. Ablation results in Tab. 7 reveal that combining 2D, 3D, and simulated data yields optimal performance. As noted in Sec. 3.3, 2D data spans indoor/outdoor scenes, enabling depth learning across scales; its removal severely degrades performance on outdoor-centric BLINK [fu2024blink]. Meanwhile, 3D data captures embodied indoor environments and mitigates the Sim2Real gap, benefiting indoor-focused CV-Bench [tong2024cambrian]. Finally, simulated data broadens spatial diversity. This tripartite data composition is thus key to effective SFT training.
Dedicated depth encoder preserves image understanding. We compare dedicated and shared image-depth encoders during SFT. In Tab. 4, the dedicated encoder better maintains image understanding under limited RGB-only data (1/20 RefSpatial QA), while the shared encoder harms general performance. Though prior work[cheng2024spatialrgpt] adopts a shared encoder, it (1) requires over twice as much RGB-only data compared to spatial-related data for co-training; (2) targets region-level depths, differing from our full-image approach.
Depth encoder improves both spatial understanding and reasoning. Recent VLMs [team2025gemini, daxberger2025mm] show that large-scale spatial training enables implicit 3D understanding (e.g., depth, distance, 3D boxes) from images alone. To assess this, we fine-tune NVILA-2B [liu2024nvila] on RefSpatial without the depth encoder, followed by continued RFT. Results indicate that depth improves single-step spatial understanding, consistent with MM-Spatial [daxberger2025mm], and yields greater gains in multi-step spatial referring. We attribute this to: (1) the need for precise coordinate prediction in spatial referring, unlike VQA’s multiple-choice; (2) cumulative reasoning across steps, amplifying the utility of depth cues.
@@ Appendix A Discussion
Depth-to-3D mapping assumption. The depth-to-3D mapping assumption is essential in our real-world evaluation, as our model predicts only 2D image-plane points, while real-world tasks typically require 3D coordinates for grasping, placement, or navigation. While depth noise and partial observations are important real-world challenges, our setting follows prior work [ji2025robobrain, yuan2024robopoint], which assumes that accurate depth-to-3D mapping is feasible given known camera intrinsics and extrinsics—sufficient for common manipulation and navigation tasks. Moreover, these challenges can be effectively mitigated via the following strategies: (1) Depth noise can be mitigated by recent advances in monocular depth estimation [piccinelli2025unidepthv2], monocular geometry prediction [wang2025moge], and stereo methods [wen2025foundationstereo]. In cases of severe noise, we employ FoundationStereo [wen2025foundationstereo] in real-world settings to mitigate this issue. (2) Partial views are mitigated in our method by leveraging pixel-level target points. Further improvement is possible by incorporating RoboRefer as a spatially-aware planner for active perception.
@@
Method CV-Bench [tong2024cambrian] $\text{BLINK}_{val}$ [fu2024blink] RoboSpatial [song2024robospatial] SAT [ray2024sat] EmbSpatial [du2024embspatial]
@@ Appendix B Implementation Details and Samples of RefSpatial Dataset
3D Embodied Video (Appx. B.2): This section outlines the 3D data selection process from CA-1M [lazarow2024cubify], discusses its limitations and mitigation strategies, and presents methods for enriched scene graph construction compared to the 2D data source. We further describe a QA generation framework that leverages detailed 3D annotations (e.g., depth maps, oriented 3D bounding boxes) to capture richer spatial relations. Finally, we detail how to generate QA pairs for the problem of “feasibility assessment for object placement in free space”.